OpenStudy (anonymous):

Using substitution to solve the system x = 2y + 3 and 4x - 5y = 9, which of the following could be the first step?

OpenStudy (anonymous):

Oh, the answer choices are: A. x - 2y = 3 4x - 5y = 9 B. 4(2y + 3) - 5y = 9 C. x + 2y = 4x - 5y D. 5x - 3y = 12 E. - 3x + 7y = -6

OpenStudy (anonymous):

What the heck? What did I do?

OpenStudy (anonymous):

I don't even know you and you are calling me ugly? Who are you?

OpenStudy (sasogeek):

meow could u kindly leave this question and please stop the hate words... thank you

OpenStudy (anonymous):

Whoa, you need to calm down meow18.. why did you just randomly come at me like that? How did you even find me on here?!

OpenStudy (sasogeek):

alain take it easy on that one :) mods are on their way :)

OpenStudy (shadowfiend):

Stop. Now.

OpenStudy (anonymous):

No I'm fine, I'm just confused as to why she is saying this stuff to me? And what is mods?

OpenStudy (shadowfiend):

Sorry about that alaina—that was aimed at meow18.

OpenStudy (sasogeek):

ok so umm lets get to your question now shall we? :)

OpenStudy (anonymous):

It's alright! Thank you everyone for helping me out though! :) You all are so wonderful and kind! <3 Thanks so much! I just was confused as to why rude things were being said to me like that randomly. Oh well, yes let's get to my question! :)

OpenStudy (shadowfiend):

And thanks for setting an example on how to engage someone who suddenly goes hostile on you. I appreciate that you didn't immediately respond with attacks :)

OpenStudy (anonymous):

I won't attack someone who has rude things to say! :) It is unnecessary! :)

OpenStudy (sasogeek):

ok so you have 2 equations and 2 variables, looks like we got some interesting stuff going on in the 2 equations though. notice that the first equation already gives u the value of x as 2y+3, by substitution, simply plug in the value of x which is (2y-3) into the second equation... solve for y and then plug in your value for y into x=2y+3, solve for x and we're done here :) difficult? I think not :) follow the instructions and you'll be happy :D

OpenStudy (sasogeek):

x=2y+3, my bad but i hope the concept is clear though, that's most important :)

OpenStudy (anonymous):

So would my answer be B. 4(2y + 3) - 5y = 9? :)

OpenStudy (anonymous):

Yes that's seems perfectly correct :)

OpenStudy (anonymous):

YAY! :) Finally I got something right! :)

OpenStudy (anonymous):

B is the Exact answer